# Persistent Topological Features in Large Language Models

### v1

@@ 3.2 The zigzag algorithm
It exploits two existing public codes that were developed for zigzag computations: Dionysus2 [54] and fastzigzag [55]. Dionysus2 is a C+++++ + library for computing persistent homology, with a specific library for zigzag persistence. In our case, it has the role of extracting the filtration $f$ and computing the $times$ array, i.e. the list of layer indices to be associated with the birth and death of the simplices across the filtration. fastzigzag allows to calculate efficiently  [^2] The algorithm performs well even for the relatively large datasets we employ for this analysis: with $10$K points embedded in a space with dimension $d=4096$, a number of neighbors for the $k_{\rm NN}$ graph of $k_{\rm NN}=10$, and a maximum homology dimension of $m=10$ on an AMD EPYC 7H12 it takes approximately $2$ hours. the persistence diagram $\rm Pers_{p}(\Phi)$ by converting the input zigzag filtration to a non-zigzag filtration of an equivalent complex with the same length, and it then converts the obtained persistence intervals back to zigzag.
3.3 Effective Persistence Image
The pairs generated within $\rm Pers_{p}(\Phi)$ are best understood by visualizing them through a persistence image, a well-known descriptor within the TDA tools. The persistence image in our case results in a grid of size $\left(2N_{\rm layers}-1\right)\times\left(2N_{\rm layers}-1\right)$, for each homology dimension $p$. Each pixel in the grid is associated with an integer value corresponding to the number of cycles appearing with that birth-death pair. Defined this way, the persistence image does not discriminate between the model and intersection layers. Their behavior is generally fairly different, and have an alternating structure between model and intersection layers. Hence, persistence images are not smooth as a function of layers. To achieve a smoother representation, we introduce effective persistence images, obtained by excluding the intersection layers from the construction. This is achieved by defining a map, similar to the approach in [56], that translates the collection of intervals from the zigzag persistence diagram of the filtration in equation 2 into intervals, where the birth and death occur only across model layers. Formally, for $b,d>0$, we obtain:
$\widehat{PI}_{p}(b/2,d/2)=PI_{p}(b,d)+PI_{p}(b-1,d)+PI_{p}(b,d-1)+PI_{p}(b-1,d-1),$ (4)
where $\widehat{PI}_{p}$ is the effective persistence image for the $p$-cycles and $b,d$ are model layers indexed by even numbers. [^3] Note that this operation does not modify the information about the model layers contained in the original ${\rm Pers}_{p}(\Phi)$, as it redefines consistently all the births and deaths. The collection of $\widehat{PI}_{p}$s taken over all $p$ contains all the information output from our zigzag algorithm, and give a useful overview of the model as a whole. On the other hand, they are not easily tractable in a statistical sense and hard to interpret. Indeed, one focus of this work is to look at the fine-grained topological structure of representation space, tracking the persistence of cycles across layers. For this purpose, we develop a suited summary of the effective persistence image in the next section.
3.4 Persistence Similarity
Given two layers $\ell_{1},\ell_{2}$, we define the persistence similarity as the fraction of $p$-cycles in $\ell_{1}$ that exist in $\ell_{2}$ as well, and have existed throughout the layers in between. Mathematically it can be expressed as
$\mathcal{S}_{p}(\ell_{1},\ell_{2})$ $=\frac{\sum_{\ell_{1}\leq M_{1},\ell_{2}>M_{2}}\widehat{PI}_{p}\left(\ell_{1},\ell_{2}\right)}{\beta_{p}(\ell_{1})}$ (5)
$M_{1}$ $=\min(\ell_{1},\ell_{2});\;\ M_{2}=\max(\ell_{1},\ell_{2})$
where $\beta_{p}(\ell)$ is the Betti number, i.e. the number of alive $p$-cycles at layer $\ell$. Given a $p$-cycle that is alive at a given layer $\ell$, we can thus define the average probability of finding it alive at any other layer as
$\bar{\mathcal{S}}_{p}(\ell)=\frac{1}{N_{\rm layers}}\sum_{\ell_{i}=1}^{N_{\rm layers}}\mathcal{S}_{p}(\ell,\ell_{i}),$ (6)
which indicates the degree of “mobility” of the system at a given layer, i.e. overall retention of cycles in each model layer. Thus, a low value of $\bar{\mathcal{S}}_{p}$ represents a phase during which internal representations are undergoing major topological changes, causing points of the dataset to change relative positions abruptly. For high values, the inverse is true, i.e. the relations among points are relatively stationary. It is worth noting that traditional measures of similarity between layers typically depend solely on their current state, namely the activation matrices on the set of data. In contrast, our method considers the trajectory from $\ell_{1}$ to $\ell_{2}$, implying that persistence similarity does not just depend on the initial and final states but also on the path between them.
4 Experiments
4.1 Models, Datasets and Benchmarks
We work with $4$ models: Llama2 [57], Llama3 [58], Mistral [59] and Pythia 6.9b [60]. These models are open-source decoder-only transformers, and they achieve high performance in the benchmarks we consider in this work. Llama2-7B, Llama3-8B, Mistral 7B, and Pythia 6.9b have 32 hidden layers, Llama2-13B has 40 hidden layers, and both Llama2-70b and Llama3-70b have 80 hidden layers.
For our purposes, the input dataset from which we take internal representations must provide a fair test of how the model processes and understands language. An extensive and accessible corpus is the Pile dataset [61], which combines $22$ datasets over a wide range of topics and structures. For computational reasons, we take the Pile-10k subset, accessible on HugginFace. [^4] https://huggingface.co/datasets/NeelNanda/pile-10k For completeness, we also consider the Standford Sentiment Treebank (SST) dataset [62]. From these datasets, each prompt is processed so that the last token is extracted at each normalization layer and the final normalization applied to the output layer.
Figure 2: Effective persistence image for the Llama 3 8B model using the SST dataset. We show $1$-cycles (Left Panel) and $2$-cycles (Right panel). The corresponding $k_{\rm NN}$ graph is constructed with $k_{\rm NN}=5$ and $k_{\rm NN}=10$, respectively. The density plot shows the amount of cycles (colorbar) for a given birth-persistence pair (x- and y-axis), where values refer to the model layer.
4.2 Zigzag persistence applied to LLM models
Effective Persistence Image.
We generate an effective persistence image for each model using the two datasets, each homology dimension up to $p=3$, and for a range of values of $k_{\rm NN}\in[1,15]$. We show an example of this effective persistent image in Figure 2 for the Llama 3 8B model for the SST dataset for $1$- and $2-$ cycles for $k_{\rm NN}=5$ and $k_{\rm NN}=15$, respectively.  [^5] The reason for using the SST dataset, instead of Pile, is that the 70B models are computationally expensive for the latter. We show that results are consistent in Appendix C. The choice of the hyperparameter $k_{\rm NN}$ is done so as to maximize the total number of cycles. The x-axis represents the layer at which a $p$-cycle is born, and the y-axis represents persistence, i.e. death layer - birth layer. The colorbar measures the amount of $p$-cycles on a given grid point. As expected, a large number of cycles are very short lived, i.e. the grid points at persistence equal unity. On the other hand, we observe that persistence is typically higher for $p$-cycles born after the first half of the model’s depth, a feature that is visually evident on the right panel of Figure 2, representing $2$-cycles, but observed across all models and dimensions, especially for $1$-cycles. A fraction of these cycles have maximal persistence, i.e. they survive until the last layer.
In computing $\widehat{PI}_{p}$s across models, dimensions and $k_{\rm NN}$ values, we observe that 00- and $3$-cycles are relatively low in number, while $1$- and $2$-cycles are higher, reaching tens of thousands of cycles per layer. This behaviour might be expected for a $k_{\rm NN}$-graph based costruction, since connections are dense even for low values of $k_{\rm NN}$, especially if points are concentrated in low dimensional regions of the representation space. We examine this behavior in detail to make sure that our construction is stable for different choices on the $k_{\rm NN}$ graph, see Appendix B for details.
The $\widehat{PI}_{p}$s from Figure 2 are suggestive of important features in the topology of internal representations, which we look at in more detail using persistence similarity. Given the prevalence of 1-cycles across various models, layers, and choices of $k_{\rm NN}$, we concentrate on these features in the following discussion.
Figure 3: Persistence similarity of $1$-cycles as defined in equation 5 for Llama 3 of two different sizes ($8$B and $70$B) and the SST dataset. For both models we fix $k_{\rm NN}=5$. A given pixel of the grid represents similarity computed between two layers. Darker regions indicate higher similarity.
Persistence similarity.
We can visualize persistence similarity, as defined in equation 5, $\mathcal{S}_{1}(\ell_{1},\ell_{2})$ as a density plot, shown in Figure 3 for Llama 3 of two different sizes ($8$B and $70$B) and the SST dataset, with corresponding plots for Pile for the $8$B shown in Appendix C. Again, we choose $k_{\rm NN}=5$ although results are similar within the $k_{\rm NN}\in[1,15]$ range. In these plots, darker regions indicate a higher fraction of $p$-cycles alive between two given layers. Note that the plot is not symmetric by definition (cfr. equation 5) meaning that at a given layer, the fraction of cycles alive at an earlier layer might be different than the ones alive at a later layer. Nevertheless, $\bar{\mathcal{S}}_{p}$ is approximately symmetric. Both models clearly show a high degree of similarity roughly midway through the depth, until before the last few layers. This is in agreement with what observed in the $\widehat{PI}_{p}$, which suggested that a $p$-cycle born after the first half of the model is likely to survive until the last layer.
Figure 4: Average Similarity as a function of model depth, computed for Llama 3 8B and varying $k_{\rm NN}$ parameter (Left Panel) and at fixed $k_{\rm NN}=5$ parameter and varying models (Right Panel).
We now compute the average similarity, i.e. the average over the column of persistence similarity (cfr. equation 6), $\bar{\mathcal{S}}_{p}$ both at fixed model and varying $k_{\rm NN}$ parameter, and at fixed $k_{\rm NN}$ parameter and varying model, as a function of the model’s depth. Results are shown in the Left and Right Panels of Figure 4, respectively. Based on the Left Panel, we choose $k_{\rm NN}=5$ as representative value for the Right Panel, given that it gives the highest values of $\bar{\mathcal{S}}_{1}$. Remarkably, $\bar{\mathcal{S}}_{1}$ peaks at the same relative depth for a wide variety of models, while the parameter $k_{\rm NN}$ only changes the normalization of the curve. This result is in agreement with previous results finding that the fraction of similar layers is approximately universal across models (e.g. [6, 7]).
4.3 Layer pruning by persistence similarity
Recently, measures of layer similarity have been used to identify layers that contribute minimally to the performance of LLMs. These layers can be pruned, and the performance re-evaluated to validate this assumption. Since persistence similarity tracks changes across layers, it can be leveraged for layer pruning by selecting layers that retain the most cycles. Consequently, we establish a pruning criterion based on average persistence similarity $\bar{\mathcal{S}}_{1}$ computed now on the Pile dataset. Specifically, we prune layers that lie within $10\%$ and $20\%$ of the maximum $\bar{\mathcal{S}}_{1}$, corresponding to conservative and aggressive pruning, respectively. Here is a schematic summary of the algorithm.
@@ 4.3 Layer pruning by persistence similarity
$\bar{\mathcal{S}}_{1},model,threshold,$
$max\leftarrow\rm{max(\bar{\mathcal{S}}_{1})}$
@@
     if $\bar{\mathcal{S}}_{1}[l]>max*threshold$ then
@@
The algorithm outputs how many and which layers have high degree of persistence similarity. We now cut those layers and measure performance using the benchmarks introduced in Section 4 and across models considered in this work.
We compare to layer pruning methods based on state of the art measures of similarity, namely [6] and [7]. Both approaches are designed to take as input the desired number of layers to prune $N_{\rm prune}$ and measure performance as $N_{\rm prune}$ grows. For a fair comparison, we feed the number of layers cut by our method as an input to the other two methods, and verify which layers they select to cut given this input, and the corresponding performance. We show a schematic diagram of the layers cut with our method (Bottom Row) and the other two methods (Upper Row) in Figure 5. Interestingly, both considered methods from [6] and [7] give the same result at fixed $N_{\rm prune}$, thus we refer to them simply as “other works”.
Figure 5: Pruned layers across models based on persistence similarity (Bottom Row) and other methods from [6, 7]. Since both these two methods give the same results, we generically call them “other works”. The number of layers pruned for all methods is defined by cutting layers that are within $10\%$ (orange) and $20\%$ (yellow) of the maximum average similarity, $\bar{\mathcal{S}}_{1}$.
The $10\%$ pruning is rather consistent across methods, with small variations. The more aggressive cut of $20\%$ generates more discrepancies, especially for Llama 3 8B, where both methods from [6] and [7] prefer to cut earlier layers.
We now show performance results in Table 1,  [^6] Results for Pyhia on MMLU tasks are not shown because the model is not designed for following the format of the tasks, as shown in [60]. where in bold we indicate the layer pruning method that has better or equal performance with respect to the other method. Despite often selecting different layers, our topology-based pruning strategy achieves generally comparable results to methods from [6] and [7], and it surpasses their result in approximately for half of the model-benchamrk combinations.

### v3

@@ Appendix B Zigzag algorithm
It exploits two existing public codes that were developed for zigzag computations: Dionysus2 [78] and fastzigzag [59]. Dionysus2 is a C+++++ + library for computing persistent homology, with a specific library for zigzag persistence. In our case, it has the role of extracting the filtration $f$ and computing the $times$ array, i.e. the list of layer indices to be associated with the birth and death of features. fastzigzag allows to calculate efficiently the persistence diagram $\rm Pers_{p}(\Phi)$ by converting the input zigzag filtration to a non-zigzag filtration of an equivalent complex with the same length, and it then converts the obtained persistence intervals back to zigzag. The computational cost of our algorithm is $\mathcal{O}(n^{2}*N_{layers})+\mathcal{O}(m^{\omega})$ where the first part is the $K_{\rm NN}$ graph creation cost for the input dataset at each layer, and the second part is the theoretical cost of FastZigZag with $\omega<2.37286$. The algorithm performs well even for the relatively large datasets we employ for this analysis: with $10$K points embedded in a space with dimension $d=4096$, a number of neighbors for the $k_{\rm NN}$ graph of $k_{\rm NN}=10$, and a maximum homology dimension of $m=10$ on an AMD EPYC 7H12 it takes approximately $2$ hours.
Appendix C A Toy Example to Visualize Zigzag Barcodes
Figure 5: Zigzag persistence analysis of calendar arithmetic tokens. Top panel: Month token representations exhibit persistent topological structure emerging at layer 5 and persisting until layer 14. The points are plotted with the first and second component of PCA, respectively X and Y axis. Bottom panel: Answer token representations show late-emerging persistent structure from layers 21-32, with the points plotted on the first and second component of PCA. For each panel, the upper half displays $k$-nearest neighbor graphs ($k=2$) for the 12 prompts across selected transformer layers, while the lower half shows the corresponding 1-dimensional persistence barcodes tracking topological feature evolution across all 33 layers. The red dotted lines in the barcode plots indicate the specific layers visualized in the upper half.
To demonstrate the zigzag persistence visualization framework, we analyze a simple calendar arithmetic task on the Mistral model. We use prompts from the form
”Let’s do some calendar math. Four months from [MONTH] is
where [MONTH] cycles through all twelve months. This example was used in [79] to analyze circular representations in language models.
We extract hidden representations from all 33 transformer layers [^14] Layer 0 is the embedding layer. and apply zigzag persistent homology with $k=2$ neighbors and maximum simplex dimension 3. The point cloud at each layer comprises 12 tokens corresponding to the 12 different month prompts. We analyze two token positions: month tokens (semantic input) and answer tokens (given by the is token). So the point cloud at each layer comprises of $12$ tokens for each prompt
Figure 5 shows the analysis for both token types. The persistence barcode for month tokens (top panel) exhibits a prominent feature from layers 5-14, while answer tokens (bottom panel) show persistent structure emerging later (layers 21-32).
The visualization demonstrates two key capabilities of zigzag barcodes:
  • Layer-wise tracking: Barcodes reveal when topological features emerge and disappear across network depth.
  • Differential patterns: Different token types exhibit distinct persistence signatures.
This toy example illustrates the framework’s ability to visualize geometric structure evolution in transformer representations.
Appendix D Combining the $k_{\rm NN}$ graph with the Vietoris-Rips complex
The k-Nearest Neighbors ($k_{\rm NN}$) complex is built by expanding the corresponding $k_{\rm NN}$ graph to a fixed dimension $m$. A key limitation of the $k_{\rm NN}$ complex is that it ranks points by proximity without considering their actual distances. As a result, once $k$ is fixed on each layer, each point is connected to its k-Nearest Neighbors, regardless of the absolute distances involved. In our setting, the number of connected components (the Betti  [^15] Betti numbers have been used in previous works [25, 30] for interpreting internal representations of neural networks. However, they describe each layer independently from the others, which is not the purpose of this work. number $\beta_{0}$) of the $k_{\rm NN}$ complexes as a function of the layers tends to be unity, i.e. the whole complex is connected, even for relatively small values of $k_{\rm NN}\gtrsim 6$. This implies that connected components contain no useful topological information on the internal representations.
To ensure consistency across layers, we select the radius $R$ in each layer such that the number of connected components, $\beta_{0}$, of the $k_{\rm NN}$ complex falls in a pre-determined range. We then compute the observables presented in this work and verify the results. For clarity, we refer to $k_{\rm NN}$ complex the construction used in the main body, and $k_{\rm NN}$-VR complexes the one presented in this section. For the sake of conciseness, we present only results for the inter-layer persistence $\bar{\mathcal{Z}}$.
In Figure 6 we show the inter-layer persistence of $1$-dimensional holes of the $k_{\rm NN}$ and the $k_{\rm NN}$-VR complexes and the 00-dimensional holes of the $k_{\rm NN}$-VR complexes computed by imposing $\beta_{0}=500\pm 100$.  [^16] We checked that results are stable as long as $\beta_{0}$ is much lower than the total number of points. We observe all three curves are qualitatively similar. This indicates the stability of the results, even when removing a considerable amount of short edges. Moreover, we observe the same behavior also on 00-dimensional holes, now that we modified the complex such that their statistics are large enough to reliably compute persistence. We argue this indicates a universal (in homology) tendency to retain relational connections among particles in the middle-late layers of the model.
Figure 6: Inter-Layer Persistence with weight $\alpha=0$ as a function of model layers computed for Llama3 8B on the SST dataset for both $k_{\rm NN}$ and $k_{\rm NN}$-VR complexes. We impose the number of connected components, $\beta_{0}=500\pm 100$ to build the $k_{\rm NN}$-VR complexes.
Appendix E Consistency of Results
E.1 Effective Persistent Images across Models
Given a fixed dataset, effective persistent images can be calculated across models and subtracted element-wise to highlight differences in how the models process the same information. We show all the possible comparisons for the $4$ models considered in this work in Figure 7, which we calculate for the SST dataset. We can observe clear patterns in differences across models, reflecting what is observed in Figure 3.
Figure 7: Element-wise difference of effective persistence image calculated for Llama 2, Llama 3, Mistral and Pythia on the SST dataset. The color bar indicates a normalized difference between $-1$ and $1$, on a logarithmic scale.
E.2 Larger Models
We verify that our topological descriptors exhibit the same qualitative results for larger models, namely Llama 2 13B, Llama 2 70B and Llama 3 70B, using the SST dataset. We show the births’ relative frequency and inter-layer persistence in Figure 8 in the left and right panels, respectively. As a representative value, we choose a weight of $\alpha=0$ for both descriptors, which gives equal weight to short- and long-lived features.
Figure 8: Births’ relative frequency (left) and inter-layer persistence (right) as a function of the models’ depth for larger models, namely Llama 2 13B, Llama 2 70B, and Llama 3 70B, compared to Llama 3 8B, computed for the SST dataset with weight $\alpha=0$.
E.3 Varying Datasets
We test our topological descriptors on $4$ different datasets, as presented in Section 4. As a reference, we consider the Llama 3 8B model and choose $\alpha=0$ as weight for both the births’ relative frequency and inter-layer persistence. We show results averaged over subsets of size $500$ points in Figure 9.
Figure 9: Births’ Relative Frequency and Inter-Layer Persistence for weight $\alpha=-0$as a function of model layers for Llama 3 8B for a range of datasets, averaged over $16$ subsets of size $500$.
We note qualitative similarity for both descriptors across all datasets, though quantitative differences can clearly be seen especially for inter-layer persistence. Interestingly, we observe that the code dataset has a slight divergent behaviour in middle layers. To investigate this further, we filter the Code dataset for 5 programming languages with different levels of verbosity (C, Java, HTML, Markdown, Python) and for each one we extract 10K prompts. We then calculate the inter-layer persistence for $\alpha=-1$ and $\alpha=2$ for these datasets so as to highlight separately short- and long-lived features. In this case, we average over $8$ subsets of size $1000$ for computational convenience. We show results in Figure 10. From these calculations, we gather that most programming languages generally exhibit a drop in the amount of short-lived features in the middle layer, the effect being more visible for verbose codes (java, C). This feature is not seen for other types of datasets. These results deserve further investigation, which we leave to future work.
Figure 10: Inter-Layer Persistence for weights $\alpha=-1$ (left panel) and $\alpha=2$ (right panel) as a function of model layers for Llama 3 8B for a range of programming languages, averaged over $8$ subsets of size $1000$.
E.4 Results as a Function of Varying Subsets
Here we show that our topological descriptors show consistent results across various subset sizes. While the variance increases for smaller subsets, descriptors computer over different subset sizes are within a standard deviation. We show these results for both the births’ relative frequency and the inter-layer persistence calculated for the Llama 3 8B model on the SST dataset in Figure 11.
Figure 11: Births’ relative frequency (left) and inter-layer persistence (right) with weight $\alpha=0$ for the Llama 3 8B model computed on the SST dataset for different subset sizes as a function of model layers.
Variance of $\bar{\mathcal{Z}}_{1}$ as a function of data points.
Given all the subsets with {100, 200, …, 1000} points, we can calculate how the variance of these subsets scales as compared to the size of the subset. As a test case, we take the Llama 3 8B model with the SST dataset and compute the inter-layer persistence at weight $\alpha=0$ over all the subsets. We then plot the variance of each subset as a function of the subset size for different layers. We choose $4$ layers so that they are roughly representative of the dynamical phases identified in the main text. We show results in Figure 12, where we overlay a fitted curve in black. Apart for the first layers, where the variance grows with the number of points, N, in later layers the relation seems to be approximately $\sigma^{2}\propto N^{3/2}$.
Figure 12: Variance of the inter-layer persistence with weight $\alpha=0$ for the Llama 3 8B model computed on the SST dataset as a function of subset sizes. We show four different layers: $3,15,23$ and $32$. The black dashed lines represent a fitting function.
E.5 Additional result across models
In this section we present supplementary evidence of the consistency of the results across models for our descriptors, effective persistence, Births’ Relative Frequency and Inter-Layer Persistence on Llama 2, Mistral, and Pythia, in Figure 13.
Figure 13: Supplementary plots for Llama 2, Mistral, and Pythia on the SST dataset. The first row displays Inter-layer persistence, the second row shows the Births’ relative frequency, and the third row presents the effective persistent images.
Appendix F A shuffling test
As a test of our topological descriptors and the phases seen in Section 4, we perform a shuffling of tokens within the prompts of the SST dataset, as a way of destroying the structure and semantic coherence of the prompts, without modifying their unigram frequency distribution (see e.g. [16] for an application of shuffling to internal representations of transformers).
In Figures 14 and 15, we show the births’ relative frequency and the inter-layer persistence for shuffled and structured prompts. For the former, we can see a clear difference in behavior on the birth of the long-lived features between the shuffled and structured cases, the peak at middle layers being higher for shuffled prompts. A reversed trend is seen for the inter-layer persistence. Overall, the frequency of births of short-lived features is not significantly affected by the shuffling, while the inter-layer persistence drops on the second half of the model’s depth. These findings deserve further investigations, which we postpone to future work.
Figure 14: Births’ relative frequency at weight $\alpha=-1$ (left) and $\alpha=2$ (right) as a function of model layers for the Llama 3 8B for shuffled and unshuffled prompts from the SST dataset.
Figure 15: Inter-layer persistence at weight $\alpha=-1$ (left) and $\alpha=2$ (right) as a function of model layers for the Llama 3 8B for shuffled and unshuffled prompts from the SST dataset.
Appendix G More on Pruning
G.1 Sliding Window on Other Benchmarks
We can test the sliding window experiments on the other two benchmarks, MMLU and Hellaswag, show in Figures 17 and 16, respectively.
Figure 16: Hellaswag 5-shot benchmark run on Llama3 8B, Llama2 7B, Mistral and Pythia. A sliding window of size 5 is applied to cut blocks every 2 layers.
Figure 17: MMLU 5-shot benchmark run on Llama3 8B, Llama2 7B and Mistral. The different benchmarks shown are done by cutting blocks of layers with a fixed size and by changing the starting point with a sliding window. Left Panel: benchmark made with a block size of 5 and sliding windows of 2, Middle Panel: benchmark made with block size of 3 and sliding windows of 1, Right Panel: with a block size of 2 and sliding window of 1.
For the case of MMLU, we zoom in on the drop in performance seen at the end of the third phase: we show the performance of the MMLU benchmark against block sizes of 5, 3, and 2 adjacent layers with sliding windows of 2, 1 and 1 for the left, middle and right panels, respectively. We see that performance is at the level of random choice during the increasing phase and it maximizes close to the maximum Inter-Layer Persistence during the plateau phase. Consistently with the Winogrande benchmark, we see a drop in performance right in correspondence with the decreasing phase. For both Llama and Mistral, the relevant layers are a few layers before the last. This finding deserves a closer investigation, which we leave for future work.
G.2 Layer pruning algorithm
Here we schematically describe the algorithm for layer pruning used to produce results presented in Table 1.
@@ G.2 Layer pruning algorithm
$\bar{\mathcal{Z}}_{1},model,threshold,$
$max\leftarrow\rm{max(\bar{\mathcal{Z}}_{1})}$
@@
     if $\bar{\mathcal{Z}}_{1}[l]>max*threshold$ then
@@
cell22 = r, cell23 = r, cell24 = r, cell32 = r, cell33 = r, cell34 = r, cell42 = r, cell43 = r, cell44 = r, cell52 = r, cell53 = r, cell54 = r, hlines, vlines, Models & $N_{\rm prune}$ Our method Other works
Llama 2 8 [20-27] [21-28]
Llama 3  9 [20-28] [20-28]
Mistral 10 [20-29] [19-28]
Pythia 11 [19-29] [18-28]
Table 2: Table with the list of models and the layers that we cut for the pruning experiment. With a given $N_{\rm prune}$ we show the layers cut with our method and for the Angular Distance and the Bi Score (Other works).
G.3 Additional benchmarks
In Table 3, we present additional benchmarks beyond those presented in the main text, using the pruning method outlined in 4.4. Results are overall consistent with what find with benchmarks in the main text.
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=3, cell15 = c=3, cell18 = c=3, cell32 = r, cell33 = r, cell34 = r, cell42 = r, cell43 = r, cell44 = r, cell52 = r, cell53 = r, cell54 = r, vlines, hline1,3-7 = -, hline2 = 2-10, Model & CMMLU Commonsense-QA WSC